# GRDD+: An Extended Greek Dialectal Dataset with Cross-Architecture Fine-tuning Evaluation

### v2

@@ Abstract
We present an extended Greek Dialectal Dataset (GRDD+)  [^1] The full code for fine-tuning and the dataset GRDD+ are available at the following anonymous link: https://drive.google.com/drive/folders/1Xwfz08S8-9ZqMGd6EaNSje33LIaSE2E5?copy.that complements the existing GRDD dataset with more data from Cretan, Cypriot, Pontic and Northern Greek, while we add six new varieties: Greco-Corsican, Griko (Southern Italian Greek), Maniot, Heptanesian, Tsakonian, and Katharevusa Greek. The result is a dataset with total size 6,374,939 words and 10 varieties. This is the first dataset with such variation and size to date. We conduct a number of fine-tuning experiments to see the effect of good quality dialectal data on a number of LLMs. We fine-tune three model architectures (Llama-3-8B, Llama-3.1-8B, Krikri-8B) and compare the results to frontier models (Claude-3.7-Sonnet, Gemini-2.5, ChatGPT-5).
@@ 1.  Introduction
Modern Greek exhibits rich dialectal variation across different geographical regions. Despite this diversity, computational resources for these dialects remain limited, constraining the study and processing of regional linguistic varieties. Meanwhile, in the current rapid advancement of Natural Language Processing (NLP), Large Language Models (LLMs) have emerged at the forefront of research and development. However, LLMs frequently struggle with dialectal variations in lower-resourced languages, e.g. in Parts of Speech (POS) tagging and dialect identification faisal2025testing. While their dialect performance can surpass zero-shot transfer, it still falls behind the fine-tuned results faisal2024dialectbench. These limitations significantly impact their ability to generate contextually appropriate responses across regional dialects.
@@ 2.  Related Work
Against this backdrop, resources for Modern Greek dialects remain scarce. Existing datasets include, among others, a small corpus for Griko anastasopoulos2018part, the Cypriot Greek version of the Multi-CAST corpus of annotated spoken texts hadjidas2015multi, and a database comprising 505 hours of recorded dialectal speech with linguistic and meta-linguistic annotations karasimos2008greed. To our knowledge, the GRDD corpus chatzikyriakidis2023grdd constitutes a first comprehensive effort to develop large-scale publicly available resources for Modern Greek dialects.
In parallel, several studies have emerged in the field of Greek computational dialectology, such as the development of Treebanks and parsers for Eastern Cretan in the framework of Universal Dependencies vakirtzian2025dialectal, the detection of Italian and Turkish loanwords in Greek dialects scherrerpsaltakichatzikyriakidis2025laonwords and computational analyses of the linguistic varieties of Cappadocian, Pharasiot, and Silliot bompolas2023computational. However, to the best of our knowledge, none of these efforts have attempted fine-tuning Large Language Models (LLMs) on Greek Dialectal Data.
@@ 3.1.  Collection Methodology
Building upon the GRDD dataset chatzikyriakidis2023grdd, which consists of four dialects of Modern Greek, specifically Cretan, Pontic, Northern Greek, and Cypriot Greek, the present work seeks to extend and enhance the resource. Specifically, we enrich the existing dialectal corpora and incorporate six additional Greek dialectal varieties, as detailed below.
@@ 3.1.1.  Greco-Corsican
In the 1670s, Greek migrants from Mani settled in Cargèse, Corsica, forming a Greek-speaking community nicholas2005history. From the 1670s to the 1960s, a span of nearly three centuries, Greek was spoken in Cargèse, in relative isolation from other Greek-speaking communities. The variety, known as Greco-Corsican, has been the subject of detailed linguistic study phardys1888yleiskarifima; blanken1951grecs; parlangeli1952dom; rexine1966vayacacos. However, linguistic assimilation progressed rapidly, and by the 1930s only about 20 speakers of Greek remained. The language ultimately became extinct with the death of its last native speaker, Justine Voglimacci, in 1976.
@@ 3.1.2.  Griko (Southern Italian Greek)
Griko is a Greek dialect spoken in Grecìa Salentina, southern Italy, and recognized as a minority language. Officially, Grecìa Salentina consists of 12 villages: Calimera, Carpignano Salentino, Castrignano dei Greci, Corigliano d’Otranto, Cutrofiano, Martano, Martignano, Melpignano, Sogliano Cavour, Soleto, Sternatia and Zollino. Griko together with Grecanico of Calabria, form the endangered Italiot Greek group salminen1999unesco. Written in the Latin alphabet and only partly intelligible with Modern Greek, Griko now has fewer than 20,000 mostly elderly speakers chatzikyriakidis2010clitics.
@@ 3.1.3.  Heptanesian
Heptanesian is a Modern Greek dialect spoken on the Ionian Islands, including Corfu, Cephalonia, Lefkada, Zante, Ithaca, Kithira, Paxi and smaller islands such as Othoni, Antipaxi, and Antikithira kontosopoulos2000nikolaou. These islands were under Venetian rule from the late 14th to the late 18th century. Heptanesian exhibits Venetian and Italian influences primarily in its vocabulary, phonology (e.g.intonation), and morphology (e.g., the noun suffix –aδa < Ven –ADA), with syntax largely unaffected ralli2012verbal. Today, Heptanesian is gradually being abandoned in favor of Standard Modern Greek (SMG).
@@ 3.1.4.  Tsakonian
Tsakonian, a highly divergent modern form of Greek, still spoken in the eastern Peloponnese, is often considered distinct enough to be classified as a separate language from the rest of Modern Greek. As the only Modern Greek dialect that is not descended from the Hellenistic Koine, Tsakonian represents the main exception among Modern Greek varieties, deriving more or less directly from the ancient Doric dialect joseph1987modern; mackridge2010modern. horrocks2014greek refers to Tsakonian as a case of extreme dialectal resilience, exempt from the fundamental sound changes that shaped Modern Greek, such as the reversal of /u/ > /i/, while the dialect also exhibits numerous features that are unusual or unique compared to other Modern Greek varieties liosis2016tsakonian.
@@ 3.1.5.  Maniot
Maniot refers to the dialect spoken in the region of Laconian Mani. According to the traditional classification proposed by hatzidakis1892einleitung, which divides Modern Greek dialects into northern and southern groups, Maniot is categorized among the southern varieties. kontosopoulos2008dialects notes that Maniot constitutes a dialect distinct from the rest of the Peloponnesian varieties. The same view appears to be supported by trudgill2003modern, which emphasize the distinctiveness of the Maniot dialect.
The linguistic systems that appear to share similarities with Maniot include SMG, Cretan trudgill2003modern, Megarian, and, of course, several other Peloponnesian dialects pantelidis2001peloponnesiakos. pantelidis2001peloponnesiakos has argued that the similarities observed between SMG and the Peloponnesian dialects result from the influence of SMG on these dialects, and not vice versa, as had previously been claimed by mackridge1994neogreek; browning1969medieval; kontosopoulos2008dialects, among others.
@@ 3.1.6.  Katharevusa Greek
Katharevusa, described as a ‘ purist ’ (literally the purifying language) variety of SMG, served as the official written language of Greece from the establishment of the modern Greek nation-state until 1976 joseph1987modern; mackridge2010modern. This language variety was the middle solution during the language controversy  [^3] The language controversy, which originated in the 1760s, re-emerged with the establishment of the modern Greek state (1830) through the debate over which variety should serve as the official language of the newly independent nation mackridge2010modern. Katharevousa emerged as a kind of compromise between adopting Ancient Greek and the spoken form of SMG as the national language horrocks2014greek., and it was mostly used in written texts mackridge2010modern (for a contrasting view that argues that conditions of diglossia were developed between Katharevousa and SMG here: joseph1987modern). Katharevusa combined elements of both Ancient and Modern Greek, retaining much of the classical vocabulary and morphology while introducing intermediate forms such as ειµεϑα ` ωε αρε ´ ” ανδ ητον ` ηε/ςηε/ιτ ωας ´. Σψνταςτιςαλλψ, ιτ ωας ςλοςερ το ΣΜΓ, υςινγ ςονςτρυςτιονς λιϰε να + φινιτε ερβ ανδ τηε νεγατιε δεν, ψετ ιτ πρεςερεδ μανψ ανςιεντ παρτιςιπιαλ ςτρυςτυρες αβςεντ φρομ τηε ςποϰεν λανγυαγεmackridge2010modern· horrocks2014greek.
@@
Τηε ΓΡΔΔ οριγιναλ ςορπυς ςομπριςες φουρ μαιν Γρεεϰ διαλεςτς: Ποντις Γρεεϰ, ῝ρεταν Γρεεϰ, ῝ψπριοτ Γρεεϰ, ανδ Νορτηερν Γρεεϰ. Ωε υςεδ τηε τερμ Ποντις Γρεεϰ το ρεφερ το τηε διαλεςτ ας ςποϰεν τοδαψ ιν μοδερν Γρεεςε, αλτηουγη α φορμ οφ Ποντις, Ρομεψϰα Ποντις, ις ςτιλλ ςποϰεν ιν ςομε ιλλαγες οφ Τραβζον ανδ ςυρρουνδινγ αρεας ιν πρεςεντ-δαψ Τυρϰεψ sitaridou2012cultural. ῝ρεταν Γρεεϰ ις ςποϰεν ον τηε ιςλανδ οφ ῝ρετε ανδ ις δεριεδ φρομ Κοινε Γρεεϰ mackridge1985moderngreek. ῝ψπριοτ Γρεεϰ ις ςποϰεν πριμαριλψ βψ Γρεεϰ ῝ψπριοτς, ας ωελλ ας βψ ςομε Τυρϰιςη ῝ψπριοτς. Τηε πρειους ερςιον οφ τηε ςορπυς ον Νορτηερν διαλεςτς ινςλυδεδ δατα ονλψ φρομ Κοζανι ανδ Γρεενα, βυτ ωε ηαε νοω εξτενδεδ ιτ το αλςο ινςλυδε Λεςβος, Σαμοτηραϰε ανδ Τηραςε ρεφλεςτινγ τηε βροαδερ ςςοπε οφ Νορτηερν διαλεςτς. Τηις ις ςομετηινγ ωορτη μεντιονινγ εεν τηουγη ωε ωιλλ νοτ διςςυςς τηε οριγιναλ ςορπυς διαλεςτς chatzikyriakidis2023grdd ιν δεταιλ.
@@
Ιν τερμς οφ τηε ινδιιδυαλ διαλεςτς, ωε νοτιςε τηατ τηε ηιγηερ ςςορες αρε γιεν το Νορτηερν ανδ ῝ρεταν ρεςπεςτιελψ, φολλοωεδ βψ ῝ψπριοτ ανδ Ποντις. Αν ιντερεςτινγ χυεςτιον ηερε ςονςερνς ωηετηερ τηις ςλινε ηας ανψτηινγ το δο ωιτη τηε διςτανςε οφ τηεςε ινδιιδυαλ διαλεςτς το τηε δομιναντ αριετψ, Μοδερν Γρεεϰ, τηατ αλλ μοδελς ηαε ατ λεαςτ ςομε ϰνοωλεδγε οφ. Ιμπρεςςιονιςτις ιντυιτιονς αβουτ τηεςε διαλεςτς διςτατε τηατ ινδεεδ Νορτηερν ανδ ῝ρεταν αρε ςλοςερ το τηε δομιναντ αριετψ, ωηιλε ῝ψπριοτ ανδ λαςτλψ Ποντις αρε φαρτηερ αωαψ. [^4] In traditional Greek dialectology, there is a distinction between idioms and dialects. Basically, idioms were varieties that were closer to the dominant but not that far away to be considered dialects, and dialects varieties that were farther away from the dominant to be considered idioms. Northern and Cretan were usually considered idioms, Pontic and Cypriot dialects Οφ ςουρςε, τηε ιςςυε οφ λινγυιςτις διςτανςε ις λαργελψ υνεξπλορεδ ιν Γρεεϰ αριετιες, βυτ ιτ ωουλδ βε ιντερεςτινγ το ςεε ωηετηερ τηεςε ρεςυλτς ηερε, ςορρελατε ωιτη ςομε νοτιον οφ διςτανςε βετωεεν τηε δομιναντ αριετψ ανδ τηε ρεςπεςτιε διαλεςτς.
Λαςτλψ, τηε ρελατιονςηιπ βετωεεν τραινινγ δατα ςιζε ανδ μοδελ περφορμανςε αςροςς τηε φουρ διαλεςτς ις ςηοων το βε χυιτε ιντριγυινγ. ῝ρεταν ις τηε φιρςτ ιν ςιζε ωιτη 9,004 εξαμπλες (44.8%) ανδ περφορμς ωελλ (2.80-3.79), ωηιλε ῝ψπριοτ ηας 6,589 εξαμπλες (32.8% οφ τηε δαταςετ) ανδ αλλ τηρεε φινε-τυνεδ μοδελς ςςορινγ αβοε 3.0, ωιτη ονε φινε-τυνεδ μοδελ (Κριϰρι-8Β ατ 2.80) φαλλινγ βελοω τηε 3.0 τηρεςηολδ. Ποντις ηας 4,190 εξαμπλες (20.8%) βυτ ςονςιςτεντλψ ςςορες λοωεςτ (2.14-2.86), ωιτη αλλ τηρεε φινε-τυνεδ μοδελς φαιλινγ το ρεαςη 3.0. Συρπριςινγλψ, Νορτηερν Γρεεϰ ωιτη ονλψ 333 εξαμπλες (1.7%), μαναγες το αςηιεε ςτρονγ ςςορες (2.84-3.86), ωιτη ονλψ ονε φινε-τυνεδ μοδελ (Λλαμα-3-8Β ατ 2.84) βελοω 3.0. Τηε φαςτ τηατ ῝ψπριοτ ις τηε ονλψ διαλεςτ ωηερε αλλ φινε-τυνεδ μοδελς ςονςιςτεντλψ μαναγε το βρεαϰ τηε 3.0 μιγητ ινδιςατε τηε βενεφιτς οφ α λαργε δαταςετ τηατ ις τηε ςομβινατιον οφ τωο διερςε ςορπορα προιδινγ βεττερ ςοεραγε. Τηε Νορτηερν ρεςυλτς ρεμαιν νοταβλε, ας τηεψ δεςπιτε ηαινγ τηε λεαςτ δατα βψ φαρ, ιτ ματςηες ῝ρεταν´ς ςονςιςτενςψ βεττερ τηαν Ποντις δοες ωιτη 12 τιμες μορε τραινινγ εξαμπλες. Τηις παττερν μιγητ βε αν ινδιςατιον οφ λινγυιςτις διςτανςε φρομ Στανδαρδ Μοδερν Γρεεϰ, δατα χυαλιτψ διφφερενςες, ορ α ςομβινατιον οφ τηε τωο.

### v3

# arXiv:2511.03772v3[cs.CL] 27 Feb 2026

@@ Abstract
We present an extended Greek Dialectal Dataset (GRDD+) that complements the existing GRDD dataset with more data from Cretan, Cypriot, Pontic and Northern Greek, while we add six new varieties: Greco-Corsican, Griko (Southern Italian Greek), Maniot, Heptanesian, Tsakonian, and Katharevusa Greek. The result is a dataset with total size 6,374,939 words and 10 varieties. This is the first dataset with such variation and size to date. We conduct a number of fine-tuning experiments to see the effect of good quality dialectal data on a number of LLMs. We fine-tune three model architectures (Llama-3-8B, Llama-3.1-8B, Krikri-8B) and compare the results to frontier models (Claude-3.7-Sonnet, Gemini-2.5, ChatGPT-5).
Keywords: Greek dialects, Large language models, Low-resource NLP
[…]
GRDD+: An Extended Greek Dialectal Dataset with Cross-Architecture Fine-tuning Evaluation
Stergios Chatzikyriakidis1† [^1] The full code for fine-tuning and the dataset GRDD+ are available at the following anonymous link: https://drive.google.com/drive/folders/1Xwfz08S8-9ZqMGd6EaNSje33LIaSE2E5?copy., Dimitris Papadakis1†footnotemark:
Sevasti-Ioanna Papaioannou2†footnotemark: , Erofili Psaltaki3†footnotemark:
1University of Crete, 2University of Athens, 3University of Turku
{stergios.chatzikyriakidis, philp0961}@uoc.gr, lt12200018@di.uoa.gr, erofili.psaltaki@utu.fi
Abstract content
22footnotetext: These authors contributed equally to this work.
@@ 1.  Introduction
Modern Greek exhibits rich dialectal variation across different geographical regions. Despite this diversity, computational resources for these dialects remain limited, constraining the study and processing of regional linguistic varieties. Meanwhile, in the current rapid advancement of Natural Language Processing (NLP), Large Language Models (LLMs) have emerged at the forefront of research and development. However, LLMs often struggle with dialectal variations in languages with less resources, e.g., in Parts of Speech (POS) tagging and dialect identification Faisal and Anastasopoulos (2025). While their dialect performance can surpass zero-shot transfer, it still falls behind the fine-tuned results Faisal et al. (2024). These limitations significantly impact their ability to generate contextually appropriate responses across regional dialects.
@@ 2.  Related Work
Against this backdrop, resources for Modern Greek dialects remain scarce. Existing datasets include, among others, a small corpus for Griko Anastasopoulos et al. (2018), the Cypriot Greek version of the Multi-CAST corpus of annotated spoken texts Hadjidas and Vollmer (2015), and a database comprising 505 hours of recorded dialectal speech with linguistic and meta-linguistic annotations Karasimos et al. (2008). To our knowledge, the GRDD corpus Chatzikyriakidis et al. (2023) constitutes a first comprehensive effort to develop large-scale publicly available resources for Modern Greek dialects.
In parallel, several studies have emerged in the field of Greek computational dialectology, such as the development of treebanks and parsers for Eastern Cretan in the framework of Universal Dependencies Vakirtzian et al. (2025), the detection of Italian and Turkish loanwords in Greek dialects Scherrer et al. (2025) and computational analyses of the linguistic varieties of Cappadocian, Pharasiot, and Silliot Bompolas (2023). However, to the best of our knowledge, none of these efforts have attempted fine-tuning Large Language Models (LLMs) on Greek Dialectal Data.
@@ 3.1.  Collection Methodology
Building upon the GRDD dataset Chatzikyriakidis et al. (2023), which consists of four dialects of Modern Greek, specifically Cretan, Pontic, Northern Greek, and Cypriot Greek, the present work seeks to extend and enhance the resource. Specifically, we enrich the existing dialectal corpora and incorporate six additional Greek dialectal varieties, as detailed below.
@@ 3.1.1.  Greco-Corsican
In the 1670s, Greek migrants from Mani settled in Cargèse, Corsica, forming a Greek-speaking community Nicholas (2005). From the 1670s to the 1960s, a span of nearly three centuries, Greek was spoken in Cargèse, in relative isolation from other Greek-speaking communities. The variety, known as Greco-Corsican, has been the subject of detailed linguistic study Φαρδύς (1888); Blanken (1951); Parlangèli (1952); Rexine (1966). However, linguistic assimilation progressed rapidly, and by the 1930s only about 20 speakers of Greek remained. The language ultimately became extinct with the death of its last native speaker, Justine Voglimacci, in 1976.
@@ 3.1.2.  Griko (Southern Italian Greek)
Griko is a Greek dialect spoken in Grecìa Salentina, southern Italy, and recognized as a minority language. Officially, Grecìa Salentina consists of 12 villages: Calimera, Carpignano Salentino, Castrignano dei Greci, Corigliano d’Otranto, Cutrofiano, Martano, Martignano, Melpignano, Sogliano Cavour, Soleto, Sternatia and Zollino. Griko together with Grecanico of Calabria, form the endangered Italiot Greek group Salminen (1999). Written in the Latin alphabet and only partly intelligible with Modern Greek, Griko now has fewer than 20,000 mostly elderly speakers Chatzikyriakidis (2010).
@@ 3.1.3.  Heptanesian
Heptanesian is a Modern Greek dialect spoken in the Ionian Islands, including Corfu, Cephalonia, Lefkada, Zante, Ithaca, Kithira, Paxi and smaller islands such as Othoni, Antipaxi, and Antikithira Kontosopoulos (2000). These islands were under Venetian rule from the late 14th to the late 18th century. Heptanesian exhibits Venetian and Italian influences primarily in its vocabulary, phonology (e.g. intonation), and morphology (e.g. the noun suffix –aδa < Ven –ADA), with syntax largely unaffected Ralli (2012). Today, Heptanesian is gradually being abandoned in favor of Standard Modern Greek (SMG).
@@ 3.1.4.  Tsakonian
Tsakonian, a highly divergent modern form of Greek, still spoken in the eastern Peloponnese, is often considered distinct enough to be classified as a separate language from the rest of Modern Greek. As the only Modern Greek dialect not descended from the Hellenistic Koine, Tsakonian represents the main exception among Modern Greek varieties, deriving more or less directly from the ancient Doric dialect Joseph et al. (1987); Mackridge (2010). Horrocks refers to Tsakonian as a case of extreme dialectal resilience, exempt from the fundamental sound changes that shaped Modern Greek, such as the reversal of /u/ > /i/, while the dialect also exhibits numerous features that are unusual or unique compared to other Modern Greek varieties Liosis (2016).
@@ 3.1.5.  Maniot
Maniot refers to the dialect spoken in the region of Laconian Mani. According to the traditional classification proposed by Hatzidakis, which divides Modern Greek dialects into northern and southern groups, Maniot is categorized among the southern varieties. Κοντοσόπουλος notes that Maniot constitutes a dialect distinct from the rest of the Peloponnesian varieties. The same view appears to be supported by Trudgill, who emphasizes the distinctiveness of the Maniot dialect.
The linguistic systems that appear to share similarities with Maniot include SMG, Cretan Trudgill (2003), Megarian, and, of course, several other Peloponnesian dialects Παντελίδης . Παντελίδης has argued that the similarities observed between SMG and the Peloponnesian dialects result from the influence of SMG on these dialects, and not vice versa, as had previously been claimed by Mackridge, 1994; Browning, 1969; Κοντοσόπουλος, 2008, among others.
@@ 3.1.6.  Katharevusa Greek
Katharevusa, described as a ‘ purist ’ (literally the purifying language) variety of SMG, served as the official written language of Greece from the establishment of the modern Greek nation-state until 1976 Joseph et al. (1987); Mackridge (2010). This language variety was the intermediate solution during the language controversy [^3] The language controversy, which originated in the 1760s, re-emerged with the establishment of the modern Greek state (1830) through the debate over which variety should serve as the official language of the newly independent nation Mackridge (2010). Katharevousa emerged as a kind of compromise between adopting Ancient Greek and the spoken form of SMG as the national language Horrocks (2014)., and it was mostly used in written texts Mackridge (2010) (for a contrasting view that argues that conditions of diglossia were developed between Katharevousa and SMG here: Joseph et al., 1987). Katharevusa combined elements of both Ancient and Modern Greek, retaining much of the classical vocabulary and morphology while introducing intermediate forms such as ειµεϑα ` ωε αρε ´ ανδ ητον ` ηε/ςηε/ιτ ωας ´. Σψνταςτιςαλλψ, ιτ ωας ςλοςερ το ΣΜΓ, υςινγ ςονςτρυςτιονς λιϰε να + φινιτε ερβ ανδ τηε νεγατιε δεν, ψετ ιτ πρεςερεδ μανψ ανςιεντ παρτιςιπιαλ ςτρυςτυρες αβςεντ φρομ τηε ςποϰεν λανγυαγε Mackridge (2010)· Horrocks (2014).
@@
Τηε ΓΡΔΔ οριγιναλ ςορπυς ςομπριςες φουρ μαιν Γρεεϰ διαλεςτς: Ποντις Γρεεϰ, ῝ρεταν Γρεεϰ, ῝ψπριοτ Γρεεϰ, ανδ Νορτηερν Γρεεϰ. Ωε υςεδ τηε τερμ Ποντις Γρεεϰ το ρεφερ το τηε διαλεςτ ςποϰεν τοδαψ ιν μοδερν Γρεεςε, αλτηουγη α φορμ οφ Ποντις, Ρομεψϰα Ποντις, ις ςτιλλ ςποϰεν ιν ςομε ιλλαγες οφ Τραβζον ανδ ςυρρουνδινγ αρεας ιν μοδερν Τυρϰεψ Sitaridou and Chatzikyriakidis (2012). ῝ρεταν Γρεεϰ ις ςποϰεν ον τηε ιςλανδ οφ ῝ρετε ανδ ις δεριεδ φρομ Κοινε Γρεεϰ Mackridge (1985). ῝ψπριοτ Γρεεϰ ις ςποϰεν πριμαριλψ βψ Γρεεϰ ῝ψπριοτς, ας ωελλ ας βψ ςομε Τυρϰιςη ῝ψπριοτς. Τηε πρειους ερςιον οφ τηε ςορπυς ον Νορτηερν διαλεςτς ινςλυδεδ δατα ονλψ φρομ Κοζανι ανδ Γρεενα, βυτ ωε ηαε νοω εξτενδεδ ιτ το αλςο ινςλυδε Λεςβος, Σαμοτηραϰε ανδ Τηραςε ρεφλεςτινγ τηε βροαδερ ςςοπε οφ Νορτηερν διαλεςτς. Τηις ις ςομετηινγ ωορτη μεντιονινγ εεν τηουγη ωε ωιλλ νοτ διςςυςς τηε οριγιναλ ςορπυς διαλεςτς Chatzikyriakidis et al. (2023) ιν δεταιλ.
@@
Ιν τερμς οφ τηε ινδιιδυαλ διαλεςτς, ωε νοτιςε τηατ τηε ηιγηερ ςςορες αρε γιεν το Νορτηερν ανδ ῝ρεταν ρεςπεςτιελψ, φολλοωεδ βψ ῝ψπριοτ ανδ Ποντις. Αν ιντερεςτινγ χυεςτιον ηερε ςονςερνς ωηετηερ τηις ςλινε ηας ανψτηινγ το δο ωιτη τηε διςτανςε οφ τηεςε ινδιιδυαλ διαλεςτς το τηε δομιναντ αριετψ, Μοδερν Γρεεϰ, τηατ αλλ μοδελς ηαε ατ λεαςτ ςομε ϰνοωλεδγε οφ. Ιμπρεςςιονιςτις ιντυιτιονς αβουτ τηεςε διαλεςτς διςτατε τηατ ινδεεδ Νορτηερν ανδ ῝ρεταν αρε ςλοςερ το τηε δομιναντ αριετψ, ωηιλε ῝ψπριοτ ανδ λαςτλψ Ποντις αρε φαρτηερ αωαψ. [^4] In traditional Greek dialectology, there is a distinction between idioms and dialects. Basically, idioms were varieties that were closer to the dominant but not that far away to be considered dialects, and dialects were varieties that were farther away from the dominant to be considered idioms. Northern and Cretan were usually considered idioms, Pontic and Cypriot dialects Οφ ςουρςε, τηε ιςςυε οφ λινγυιςτις διςτανςε ις λαργελψ υνεξπλορεδ ιν Γρεεϰ αριετιες, βυτ ιτ ωουλδ βε ιντερεςτινγ το ςεε ωηετηερ τηεςε ρεςυλτς ηερε, ςορρελατε ωιτη ςομε νοτιον οφ διςτανςε βετωεεν τηε δομιναντ αριετψ ανδ τηε ρεςπεςτιε διαλεςτς.
Λαςτλψ, τηε ρελατιονςηιπ βετωεεν τραινινγ δατα ςιζε ανδ μοδελ περφορμανςε αςροςς τηε φουρ διαλεςτς ις ςηοων το βε χυιτε ιντριγυινγ. ῝ρεταν ις τηε φιρςτ ιν ςιζε ωιτη 9,004 εξαμπλες (44.8%) ανδ περφορμς ωελλ (2.80-3.79), ωηιλε ῝ψπριοτ ηας 6,589 εξαμπλες (32.8% οφ τηε δαταςετ) ανδ αλλ τηρεε φινε-τυνεδ μοδελς ςςορινγ αβοε 3.0, ωιτη ονε φινε-τυνεδ μοδελ (Κριϰρι-8Β ατ 2.80) φαλλινγ βελοω τηε 3.0 τηρεςηολδ. Ποντις ηας 4,190 εξαμπλες (20.8%) βυτ ςονςιςτεντλψ ςςορες λοωεςτ (2.14-2.86), ωιτη αλλ τηρεε φινε-τυνεδ μοδελς φαιλινγ το ρεαςη 3.0. Συρπριςινγλψ, Νορτηερν Γρεεϰ ωιτη ονλψ 333 εξαμπλες (1.7%), μαναγες το αςηιεε ςτρονγ ςςορες (2.84-3.86), ωιτη ονλψ ονε φινε-τυνεδ μοδελ (Λλαμα-3-8Β ατ 2.84) βελοω 3.0. Τηε φαςτ τηατ ῝ψπριοτ ις τηε ονλψ διαλεςτ ωηερε αλλ φινε-τυνεδ μοδελς ςονςιςτεντλψ μαναγε το βρεαϰ τηε 3.0 μιγητ ινδιςατε τηε βενεφιτς οφ α λαργε δαταςετ τηατ ις τηε ςομβινατιον οφ τωο διερςε ςορπορα, προιδινγ βεττερ ςοεραγε. Τηε Νορτηερν ρεςυλτς ρεμαιν νοταβλε, ας τηεψ δεςπιτε ηαινγ τηε λεαςτ δατα βψ φαρ, ιτ ματςηες ῝ρεταν´ς ςονςιςτενςψ βεττερ τηαν Ποντις δοες, ωιτη 12 τιμες μορε τραινινγ εξαμπλες. Τηις παττερν μιγητ βε αν ινδιςατιον οφ λινγυιςτις διςτανςε φρομ Στανδαρδ Μοδερν Γρεεϰ, δατα χυαλιτψ διφφερενςες, ορ α ςομβινατιον οφ τηε τωο.